OpenStudy (anonymous):

2. In a set of 400 values, the mean is 121 and the standard deviation is 14. Estimate how many values fall in between each interval a) 107 – 135= b)93 - 149= c) 79 - 163=

OpenStudy (vaboro):

Normal distribution?

OpenStudy (anonymous):

yes standard deviation

OpenStudy (vaboro):

If this is normal, then 68.4% of values fall in between plus/minus sigma (within two sigmas), 95.4% - within plus/minus two sigmas (within four sigmas) and 99.6% - within six sigmas.

jimthompson5910 (jim_thompson5910):

a) Both 107 and 135 are 1 standard deviation away from the mean. So by the empirical rule, about 68% of the distribution is in this interval. So 68% of 400 is 0.68*400 = 272 which means that there are roughly 272 values in this interval. ------------------------------------------------ b) Here 95% of the distribution lives because we're now 2 standard deviations away from the mean. So 0.95*400 = 380 values are in this range ------------------------------------------------ c) We're now 3 standard deviations away from the mean. So 99.7% of the values are found here. 0.997 * 400 = 398.8 So roughly 399 values are found here.

OpenStudy (vaboro):

http://en.wikipedia.org/wiki/Normal_distribution note distribution histogram on the page